Required information The following information applies to the questions displayed below Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units): Sales Variable expenses Contribution margin Fixed expenses $ 15,000 9,000 6,000 3,120 $ 2,880 Net operating income 9. What is the break-even point in dollar sales?

Required information The following information applies to the questions displayed below. Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units): Sales Variable expenses Contribution margin Pixed expenses Net operating income s 15,000 6, 000 $ 2,880 11. What is the margin of safety in dollars? What is the margin of safety percentage? Margin of safety in dollars Margin of safety percentage

Required informatiorn The following information applies to the questions displayed below.) Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units): Sales Variable expenses Contribution margin Fixed expenses Net operating income $ 15,000 9,000 6,000 3,120 $ 2,880 14. Assume that the amounts of the companys total variable expenses and total fixed expenses were reversed. in other words, assume that the total variable expenses are $3,120 and the total fixed expenses are $9,000. Under this scenario and assuming that total sales remain the same, what is the degree of operating leverage? (Round your answer to 2 decimal places.) of

Required information The following information applies to the questions displayed below.) Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units): Sales Variable expenses Contribution margin Fixed expenses Net operating income $ 15,000 9.000 6,000 3, 120 $ 2,880 15. Assume that the amounts of the companys total variable expenses and total fixed expenses were reversed. In other words, assume that the total variable expenses are $3,120 and the total fixed expenses are $9,000. Given this scenario and assuming that total sales remain the same. Using the degree of calculated operating leverage, what is the estimated percent increase in net operating income of a 5% increase in sales? (Round your intermediate calculations and final answer to 2 decimal places.)
Answer #1

9. Break even sales= Total fixed/(selling price-variable cost)

   Break even sales= 3120/(15-9)=520 units

         Break even in $ sales=520*15= $7,800

11. Margin of safety= Actual sales- Break even point = 15,000-7,800 = 7,200

Margin of safety percentage = Margin of safety / Actual sale = 7,200/15000= 48%

14. Degree of Operating Leverage = contribution margin/ operating income

   DOL = 11,880/2,880 = 4.12

15.percentage increase in Net Operation Income:

            4.12 = X / 5

   X = 20.6

        Increase in Net Operating Income is 20.6